OpenStudy (anonymous):

A linear function goes through the points (4,-3) and (9,7). find point-slope form, slope-intercept form, and standard form

OpenStudy (anonymous):

So first the point slop form. (y-y1)=m(x-x1) You need to find the slope of the two points. \[\frac{ -3-7 }{ 4-9}= 2\]

OpenStudy (anonymous):

Then plug in the information into the points. Point-slope, as the name suggests, rquires just a slope and one point. Because we found the slope, we can just pick one point and plug that in. (y+3)=2(x-4)

OpenStudy (anonymous):

For slope-intercept (y=mx+b) you just need to expand the point-slope form. (y+3)=2(x-4) -> y= 2x-8-3 -> y=2x-11

OpenStudy (anonymous):

Standard form is when all the variables are on the left side of the equation and the constant is on the right. Just use the point-slope and manipulate the equation (y+3)=2(x-4) -> y+3-2x=-8 -> -2x + y =-11

OpenStudy (anonymous):

so what is the point slope form im confused on the answer

OpenStudy (anonymous):

@ArkGoLucky

OpenStudy (anonymous):

The point slope for is the first answer I gave

OpenStudy (anonymous):

*form

OpenStudy (anonymous):

After the slope

OpenStudy (anonymous):

(y+3)=2(x-4) A line can defined by a point and slope so this equation gives the slope, in this case 2, and a point, in this case (4,-3)